OpenStudy (anonymous):

4x^3 - 3x^2 + x + 1 divided by x^2 + 3 how to divide using long division form

OpenStudy (anonymous):

first 4x times: \[4x^{3}-3x^{2}+x+1-4x^{3}-12x=-3x^{2}-11x+1\] -3 times: \[-3x^{2}-11x+1+3x^{2}+9=-11x+10\] so we get 4x-3 with remainder -11x+10

OpenStudy (anonymous):

4x - 3 ___________________________ x^2 + 3 | 4x^3 - 3x^2 + x + 1 -------- -4x^3 - 12x ----------------------- - 3x^2 - 11x + 1 3x^2 + 9 ----------------- - 11x + 10 so quotient = 4x -3 remainder = -11x + 10